OpenStudy (anonymous):

Solve x + 3y = 7 if x and y must be whole numbers. {(0, 5), (1, 3), (2, 1)} {(2, 1), (1, 4), (0, 7)} {(1, 2), (4, 1), (7, 0)} {(0, 5), (1, 4), (2, 1)}

OpenStudy (anonymous):

You can plug in the values in and see which of them is correct. A coordinate is (x,y), so you can put the coordinates into the equation and see which set is correct. If you start with the first one: We'll put in (0,5): 0+3(5)=7 15=7 is not true so A can't be the answer. Next we'll go to B and use (2,1): 2+3(1)=7 5=7 is not true so B can't be the answer. Now we'll go to C and use (1,2): 1+3(2)=7 7=7 this is correct so we'll use the second pair of coordinates in C, (4,1): 4+3(1)=7 7=7 This is correct so we'll now use the last pair in the set (7,0): 7+3(0)=7 7=7 Since all three of the coordinate pairs worked in C, C would be your answer :)

OpenStudy (anonymous):

Thanks :)